# A Statistician’s Overview of Mechanistic-Informed Modeling

### v1

@@ Abstract
The recent success of deep neural network models with physical constraints (so-called, Physics-Informed Neural Networks, PINNs) has led to renewed interest in the incorporation of mechanistic information in predictive models. Statisticians and others have long been interested in this problem, which has led to several practical and innovative solutions dating back decades. In this overview, we focus on the problem of data-driven prediction and inference of dynamic spatio-temporal processes that include mechanistic information, such as would be available from partial differential equations, with a strong focus on the quantification of uncertainty associated with data, process, and parameters. We give a brief review of several paradigms and focus our attention on Bayesian implementations given they naturally accommodate uncertainty quantification. We then specify a general Bayesian hierarchical modeling framework for spatio-temporal data that can accommodate these mechanistic-informed process approaches. The advantage of this framework is its flexibility and generalizability. We illustrate the methodology via a simulation study in which a nonlinear Burgers’ equation PDE is embedded via a Bayesian PINN.
@@ 1 Introduction
As Galileo supposedly noted, “Mathematics is the language in which God has written the universe.” Indeed, much of our understanding of the world around us, including physical laws and principles, is codified in terms of mechanistic formulas, such as partial and ordinary differential equations. From a statistical perspective, one can attempt to infer mechanistic equations that describe the system under study. Moreover, one can use a priori knowledge of mechanistic equations describing a physical process of interest, along with data, for inference (e.g., learning the parameters that control the specified equations or initial conditions) or prediction (e.g., forecasting/interpolating).
@@
When data are involved, one inherently has uncertainty when building models and optimization methods become an integral part of understanding the world around us. The classic example is when Carl Friedrich Gauss used the method of least squares in the early 1800s to successfully calculate the orbits of celestial bodies; importantly, Gauss used Kepler’s laws of celestial mechanics as constraints to his least squares solution (Gauss,, 1809). Indeed, statisticians have long been interested in using proposed mathematical relationships in the presence of observations to motivate statistical models. For example, as described in Wikle & Hooten, (2010), in 1927, Yule used pendulum motion differential equations to suggest an autoregressive time series for sunspot observations and Hotelling used approximations of differential equations to develop models for population growth in the United States. The integration of mechanistic information in statistical models is now a fundamental component of several sub-disciplines of statistics; e.g., statistical meteorology (Berliner et al.,, 1999), ecological statistics (Hefley et al.,, 2017), epidemiology (Leach et al.,, 2020), glaciology (Gopalan et al.,, 2018), statistical oceanography (Herbei et al.,, 2008; Milliff et al.,, 2011), pharmacokinetic and pharmacodynamic analysis in the development of pharmaceuticals (e.g., Gabrielsson & Weiner,, 2001), to name a few.
@@
With the advent of deep learning, the use of mechanistic information combined with deep neural network models has seen an exponential increase in applications across many disciplines. Such approaches, often known as Physics Informed Neural Networks (PINNs, Raissi et al., 2017a, ), are typically based on augmenting the neural network loss (e.g., mean squared error) with loss terms corresponding to the physical equations and boundary/initial conditions (e.g., Karniadakis et al.,, 2021). Although this is novel in the context of neural network models, augmenting loss functions to include mechanistic information when fitting models to data is not new. For example, Sasaki pioneered the approach in a variational solution to the data assimilation problem in meteorology beginning in the 1950s (e.g., Sasaki,, 1958, 1970). Around the same time, control engineers were combining (linearized) dynamical equations within a filter that updated in the presence of new observations (e.g., Kalman,, 1960). Additional approaches developed in engineering and statistics to “fuse” data, random processes, and deterministic models were developed starting in the 1990s and continue to this day (e.g., see overviews in Berliner,, 2003; Cockayne et al.,, 2019; Meng et al.,, 2020).
@@
From a statistical perspective, the goal is not to develop a “solver” for a known mechanistic set of equations as is often of interest in the PINN literature (Raissi et al.,, 2019), but rather to recognize, in the G. Box sense, that all our mechanistic models are in some way “wrong” when applied to real world scenarios, and that we need to account for the uncertainty in the model specification, controlling parameters, and data. In this sense, the Bayesian hierarchical modeling (BHM) approach has long been seen as the natural framework to account for all of these sources of uncertainty in physics-motivated statistical modeling (e.g., see Berliner et al.,, 1999; Cressie & Wikle,, 2011; Hennig et al.,, 2015). The Bayesian view has also increasingly been shared in the PINN community (e.g., Yang et al.,, 2021; Meng et al.,, 2022; Huang et al.,, 2023; Psaros et al.,, 2023). This has been an important development from a statistical perspective. Yet, some components of the BHM modeling paradigm are often not explicitly represented in these implementations, and there is still very little consideration of the importance of specifying dependent random effects, which has been shown to be very useful in statistical mechanistic modeling of spatio-temporal processes (e.g., Wikle et al.,, 2001; Cressie & Wikle,, 2011; Chkrebtii et al.,, 2016).
@@
Our goal in this paper is to present a statistically-minded overview of mechanistically-informed modeling in the presence of data and uncertainty. Note, we prefer the term “mechanistically-informed” to “physics-informed” because many mechanistic models for non-physical processes (e.g., invasive species in ecology, Wikle, (2003)) also follow this paradigm and it can be misleading to use the term “physics” for such processes. We will approach this from a general BHM perspective, incorporating both neural and statistical surrogate models for the mechanistic process (in our case, PDEs). We first provide a brief historical overview of how mechanistic process information has been combined with data in various modeling paradigms. Throughout, we discuss issues of interest to statisticians, such as uncertainty in data, parameters, and process knowledge, as well as computational tractability. We then show how the PINN perspective fits naturally into a general BHM framework.
@@
Section 2 discusses the general problem at hand and introduces the relevant notation. Then, in Section 3 we present a brief review of mechanistically-informed modeling from the perspective of variational data assimilation and optimal interpolation, as well as statistical approaches, with a focus on Bayesian and hierarchical Bayesian implementations to account for uncertainty. Section 4 describes PINNs and Bayesian-PINNs (B-PINNs) as well as connections to the aforementioned earlier literature. Section 5 discusses a general BHM framework that can accommodate mechanistically-driven process models (e.g., physical-statistical models and B-PINNs). We then demonstrate in Section 6 this general approach with a B-PINN corresponding to a Burgers’ equation mechanistic PDE via a simulation study. We conclude with a brief overview and discussion of challenges in Section 7.
@@ 2 Problem Outline
Our interest in mechanistically-informed models here concerns problems common to spatio-temporal statistics where one has potentially incomplete and noisy observations of a univariate or multivariate process at some locations in space and time, and seeks to: (a) determine an initial condition (i.e., the true process values at specified spatial locations at some initial time), (b) predict (interpolate) incomplete observations at specified locations in space and time, or (c) forecast at specified spatial locations at times beyond which we have observations. We generally assume that we have some a priori knowledge of the underlying spatio-temporal dynamics of the process, yet our knowledge is incomplete. That is, we may not know the controlling parameters of the mechanistic model and/or the mechanistic model is an incomplete representation of the process. (An incomplete representation of dynamics could be due to many circumstances – for instance, due to approximations to the underlying partial differential equations, or in other instances due to missing mechanistic understanding.) Next, we outline the notation to be used in this exposition.
@@ 2.1 Process and Data Notation
Assume we have observations of elements of the process corresponding to some or all of the times in $\mathcal{T}$ at $m_{t}$ spatial locations $\bm{r}_{j}\in\mathcal{D}_{m}\subseteq\mathcal{D}$, which we denote by the $m_{t}$-vector $\bm{z}_{t}$. Some or all of these observation locations may coincide with the spatial locations in $\mathcal{S}$, but may not, as discussed below. Note, it may be the case that $m_{t}=0$ for some times (i.e., there are no observations for certain times).
@@ 2.2 Mechanistic Model Notation
Here we focus our attention on mechanistic models that take the form of a time-dependent partial differential equation (PDE), but the ideas presented here can be easily adapted to other forms of mechanistic models (e.g., ordinary differential equations, or agent-based models). Assume we have a PDE that we believe is mechanistically informative to our latent process, $\bm{u}_{t}(\bm{s})$. For example, we may believe that our latent process shows evidence of advection and diffusive dynamics. We denote this
$\frac{\partial\tilde{\bm{u}}_{t}(\bm{s})}{\partial{t}}=\mathcal{M}_{u}(\tilde{\bm{u}}_{t}(\bm{s}),\bm{\theta}_{m}),$ (1)
where $\tilde{\bm{u}}_{t}(\bm{s})$ corresponds to the $K$-dimensional mechanistic process defined at $t\in[0,T+\tau]$ and $\bm{s}\in\mathcal{D}$ and $\mathcal{M}_{u}$ is an operator controlled by parameters $\bm{\theta}_{m}$. Note, we are deliberately making a distinction between the true latent process of interest, $\bm{u}_{t}(\bm{s})$ (no tilde), and the mechanistic model representation, given by $\tilde{\bm{u}}_{t}(\bm{s})$ this distinction will be made more clear when we discuss models. Assume that the PDE has associated initial and boundary conditions, respectively,
$\tilde{\bm{u}}_{0}(\bm{s})$ $=\bm{g}_{I}(\bm{s}),\;\;\bm{s}\in\Omega,$
$\bm{B}(\tilde{\bm{u}}_{t}(\bm{s}))$ $=0,\;\;t\in[0,T+\tau],\bm{s}\in\mathcal{D}_{B},$
@@
Our goal is to perform inference on $\bm{u}_{t}(\bm{s})$ at the $\mathcal{S}$ spatial locations and at times $t\in\mathcal{T}$ while taking into account the information present in $\mathcal{M}_{u}$ and given the data $\bm{Z}\equiv\{\bm{z}_{1},\ldots,\bm{z}_{T}\}$. We can denote this as a posterior distribution (with some abuse of notation) as $[\bm{u}_{t}:t\in\mathcal{T}|\bm{Z},\mathcal{M}_{u}]$, where we use the bracket notation ([][\;\;][ ]) to represent a probability distribution, as is common in the Bayesian hierarchical modeling literature (e.g., Cressie & Wikle,, 2011); that is, $[a|b]$ corresponds to the distribution of $a$ given (conditional upon) $b$, $[b]$ corresponds to the marginal distribution of $b$, and $[a,b]$ corresponds to the joint distribution of $a$ and $b$.
In practice, we may be interested in learning the initial condition ($\bm{u}_{0}$), interpolation at some time $t$ ($\bm{u}_{t}$), or forecasting ($\bm{u}_{T+\tau}$). This is generally termed the “forward problem” where $\mathcal{D},\mathcal{D}_{B},\bm{g}_{I}(\cdot),\bm{B}(\cdot)$, and $\bm{\theta}_{m}$ are known and one seeks to find $\bm{u}_{0}(\bm{s})$ and $[\bm{u}_{t}]_{t\in[0,T+\tau]}$ given the dynamics suggested by $\mathcal{M}_{u}$. We may also be interested in learning parameters associated with the mechanistic PDE ($\bm{\theta}_{m}$) and various parameters associated with the underlying statistical or neural models used to represent the PDE (see below). This is generally termed the “inverse problem,” where we may have partial observations of the process in the domain of interest. In practice, we are often interested in learning about both $[\bm{u}_{t}]_{t\in[0,T+\tau]}$ on $\mathcal{D}$ and $\bm{u}_{0}(\bm{s})$ on $\mathcal{D}_{B}$ as well as the $\bm{\theta}_{m}$ that satisfies $\mathcal{M}_{u}$. This is a major focus of the methodology presented below.
@@
$\tilde{\bm{u}}_{t+1}=\mathcal{M}(\tilde{\bm{u}}_{t};\bm{\theta}_{m}),$ (2)
but recognize that there are PDE solvers that provide $\tilde{\bm{u}}_{t}$ for any and all times in $[0,T]$ that are not necessarily sequential in time (e.g., implicit solvers). Note, we drop the subscript $u$ on $\mathcal{M}$ to distinguish this numerical operator from the theoretical operator given in (1).
@@ 3 Historical Perspective
3.1 Variational Data Assimilation Approach
@@
It was known shortly after L.F. Richardson’s failed attempt at the first numerical weather forecast in the early 1920s (Richardson,, 1922) that the numerical solution of PDEs with initial conditions given by data would quickly become unstable due to lack of consistency with the governing equations. Thus, by the time that computers were actually capable of generating a realistic forecast in real time, it was known that initial condition fields of weather variables had to be smoothed in a way that respected dynamical constraints (Charney et al.,, 1950; Phillips,, 1956). Two main approaches were developed in the 1950s and 1960s to do this, one deterministic based on the calculus of variations and one stochastic, based on what would come to be known as Gaussian processes (GPs). These approaches are essentially equivalent when one takes a Bayesian inferential perspective (e.g., Wikle & Berliner,, 2007). In this section, we briefly discuss these approaches as they provide the foundation for mechanistically-informed models for spatio-temporal processes in statistics.
@@
The initial idea in data assimilation that one could simultaneously minimize an objective function containing both the sum of squared differences between the observations and the “true” initial field as well as a constraint that incorporated the PDE dynamics was presented in Sasaki, (1958). Although he formulated this in continuous space, practical uses of the methodology considered discrete space (e.g., Thompson,, 1969; Sasaki,, 1970; Stephens,, 1970).
@@
Following the notation in Section 2.1, say we have $m_{t}$-dimensional data vectors, $\bm{z}_{t}$, and numerical PDE solutions of the latent (state) process at time $t$ denoted by $\tilde{\bm{u}}_{t}$ given by (2). Here, without loss of generality, we assume $K=1$ (a univariate process) for notational simplicity. Critically, for deterministic systems, if one knows the initial state $\tilde{\bm{u}}_{0}$ then (2) determines the sequence of future states, $\{\tilde{\bm{u}}_{1},\tilde{\bm{u}}_{2},\ldots,\tilde{\bm{u}}_{t},\ldots\}$. In Sasaki’s variational approach, one then breaks the objective function into two components $J_{0}(\tilde{\bm{u}}_{0})$ and $J_{1}(\tilde{\bm{u}}_{0})$, given as follows (e.g., see Lewis & Lakshmivarahan,, 2008). First, let
$J_{0}(\tilde{\bm{u}}_{0})=\frac{1}{2}\sum_{t=1}^{T}(\bm{z}_{t}-\bm{h}(\tilde{\bm{u}}_{t}))^{\prime}\bm{R}_{t}^{-1}(\bm{z}_{t}-\bm{h}(\tilde{\bm{u}}_{t})),$ (3)
@@
where $\bm{h}(\cdot)$ is a (known) mapping function that takes the $n$-dimensional deterministic model state vector $\tilde{\bm{u}}_{t}$ and maps to the $m_{t}$-dimensional observation vector at time $t$ (e.g., an interpolator if data are not coincident with model locations, or perhaps the identity if they are; this function can also take into account nonlinear transformations, see Daley, (1993)). In addition, $\bm{R}_{t}$ is the time-varying $m_{t}\times m_{t}$ observation error covariance matrix (representing measurement error and other sources of observational error), also typically assumed known. The physical constraint is then given by:
$J_{1}(\bm{\lambda},\tilde{\bm{u}}_{0})=\sum_{t=1}^{T}\bm{\lambda}_{t}^{\prime}(\tilde{\bm{u}}_{t+1}-\mathcal{M}(\tilde{\bm{u}}_{t})),$ (4)
where $\bm{\lambda}_{t}$ are $n\times 1$ vectors of Lagrange multipliers and $\bm{\lambda}\equiv\{\bm{\lambda}_{1},\ldots,\bm{\lambda}_{T}\}$. Thus, the optimal initial state vector is found by minimizing the loss function: $L(\bm{\lambda},\tilde{\bm{u}}_{0})=J_{0}(\tilde{\bm{u}}_{0})+J_{1}(\bm{\lambda},\tilde{\bm{u}}_{0})$, where we note that $J_{1}$ is implicitly a function of the initial state, $\tilde{\bm{u}}_{0}$, through the dynamic model. Indeed, the constraints here come through the dynamic model, $\mathcal{M}(\cdot)$. Given that this model is deterministic and assumed to be “perfect”, this is considered a “strong” constraint –i.e., the solution must follow the deterministic model (Sasaki,, 1970).

### v2

# A Statistician’s Overview of Physics-Informed Neural Networks for Spatio-Temporal Data

@@ Abstract
The recent success of deep neural network models with physical constraints (so-called, Physics-Informed Neural Networks, PINNs) has led to renewed interest in the incorporation of mechanistic information in predictive models. Statisticians and others have long been interested in this problem, which has led to several practical and innovative solutions dating back decades. In this overview, we focus on the problem of data-driven prediction and inference of dynamic spatio-temporal processes that include mechanistic information, such as would be available from partial differential equations, with a strong focus on the quantification of uncertainty associated with data, process, and parameters. We give a brief review of several paradigms and focus our attention on Bayesian implementations given they naturally accommodate uncertainty quantification. We then show that it is straight-forward to include the Bayesian PINN (B-PINN) within the Bayesian hierarchical model (BHM) framework that has long been considered for modeling dynamic spatio-temporal processes. Such a BHM-PINN is illustrated via a simulation study in which a latent nonlinear Burgers’ equation PDE governs the dynamics of Poisson distributed spatio-temporal data.
@@ 1 Introduction
Paraphrasing Galileo, the laws of nature are written in the language of mathematics. Indeed, much of our understanding of the world around us, including physical laws and principles, is codified in terms of mechanistic formulas, such as partial and ordinary differential equations. From a statistical perspective, one can attempt to infer mechanistic equations that describe the system under study. Moreover, one can use a priori knowledge of mechanistic equations describing a physical process of interest, along with data, for inference (e.g., learning the parameters that control the specified equations or initial conditions) or prediction (e.g., forecasting/interpolating).
@@
When data are involved, one inherently has uncertainty when building models and optimization methods become an integral part of understanding the world around us. The classic example is when Carl Friedrich Gauss used the method of least squares in the early 1800s to successfully calculate the orbits of celestial bodies; importantly, Gauss used Kepler’s laws of celestial mechanics as constraints to his least squares solution (gauss1809theoria). Indeed, statisticians have long been interested in using proposed mathematical relationships in the presence of observations to motivate statistical models. For example, as described in wikle2010, in 1927, Yule used pendulum motion differential equations to suggest an autoregressive time series for sunspot observations and Hotelling used approximations of differential equations to develop models for population growth in the United States. The integration of mechanistic information in statistical models is now a fundamental component of several sub-disciplines of statistics; e.g., statistical meteorology (berliner1999bayesian), ecological statistics (hefley2017mechanism), epidemiology (leach2020linking), glaciology (gopalan2018bayesian), statistical oceanography (herbei2008gyres; milliff2011ocean), pharmacokinetic and pharmacodynamic analysis in the development of pharmaceuticals (e.g., gabrielsson2001pharmacokinetic), to name a few.
@@
With the advent of deep learning, the use of mechanistic information combined with deep neural network models has seen an exponential increase in applications across many disciplines. Such approaches, often known as Physics Informed Neural Networks (PINNs, Raissi2017), are typically based on augmenting the neural network loss (e.g., mean squared error) with loss terms corresponding to the physical equations and boundary/initial conditions (e.g., Karniadakis). Although this is novel in the context of neural network models, augmenting loss functions to include mechanistic information when fitting models to data is not new. For example, Sasaki pioneered the approach in a variational solution to the data assimilation problem in meteorology beginning in the 1950s (e.g., sasaki1958objective; sasaki1970some). Around the same time, control engineers were combining (linearized) dynamical equations within a filter that updated in the presence of new observations (e.g., kalman1960new). Additional approaches developed in engineering and statistics to “fuse” data, random processes, and deterministic models were developed starting in the 1990s and continue to this day (e.g., see overviews in berliner2003physical; cockayne2019bayesian; meng2020survey).
@@
From a statistical perspective, the goal is not to develop a “solver” for a known mechanistic set of equations as is often of interest in the PINN literature (Raissi2019), but rather to recognize, in the G. Box sense, that all our mechanistic models are in some way “wrong” when applied to real world scenarios, and that we need to account for the uncertainty in the model specification, parameters, and data. In this sense, the Bayesian hierarchical modeling (BHM) approach has long been seen as the natural framework to account for all of these sources of uncertainty in physics-motivated statistical modeling (e.g., see berliner1999bayesian; cressie2011statistics; hennig2015probabilistic). The Bayesian view has also increasingly been shared in the PINN community (e.g., yang2021b; meng2022learning; huang2023posterior; psaros2023uncertainty). This has been an important development from a statistical perspective. Yet, some components of the BHM modeling paradigm are often not explicitly represented in these implementations, and there is still very little consideration of the importance of specifying dependent random effects, which has been shown to be very useful in statistical mechanistic modeling of spatio-temporal processes (e.g., wikle2001spatiotemporal; cressie2011statistics; chkrebtii2016bayesian).
@@
Our goal in this paper is to present a statistically-minded overview of mechanistically-informed modeling in the presence of data and uncertainty to provide context in our discussion of PINNs. Note, although we generally use the PINN acronym due to its ubiquitous use in the literature, we prefer the term “mechanistically-informed” to “physics-informed” because many mechanistic models for non-physical processes (e.g., invasive species in ecology, wikle2003hierarchical) also follow this paradigm and it can be misleading to use the term “physics” for such processes. We will illustrate how it is natural to consider a PINN within a traditional BHM as is commonly considered for dynamic spatio-temporal statistical models (e.g., see cressie2011statistics); we call this a BHM-PINN. We first provide a brief historical overview of how mechanistic process information has been combined with data in various modeling paradigms. Throughout, we discuss issues of interest to statisticians, such as uncertainty in data, parameters, and process knowledge, as well as computational tractability.
@@
Section 2 discusses the general problem at hand and introduces the relevant notation. Then, in Section 3 we present a brief review of mechanistically-informed modeling from the perspective of variational data assimilation and optimal interpolation, as well as statistical approaches, with a focus on Bayesian and hierarchical Bayesian implementations to account for uncertainty. Section 4 describes PINNs and Bayesian-PINNs (B-PINNs) as well as connections to the aforementioned earlier literature. Section 5 discusses a how it is natural to include a B-PINN in a traditional BHM framework for dynamic spatio-temporal models, which we call a BHM-PINN. In Section 6, we apply the BHM-PINN framework to Poisson distributed spatio-temporal data where the latent process corresponds to a mechanistic PDE parameterized using Burgers’ equation. We conclude with a brief overview and discussion of challenges in Section 7.
@@ 2 Problem Outline
Our interest in mechanistically-informed models here concerns problems common to spatio-temporal statistics where one has potentially incomplete and noisy observations of a univariate or multivariate process at some locations in space and time, and seeks to: (a) determine an initial condition (i.e., the true process values at specified spatial locations at some initial time), (b) predict (interpolate) incomplete observations at specified locations in space and time, or (c) forecast at specified spatial locations at times beyond which we have observations. We generally assume that we have some a priori knowledge of the underlying spatio-temporal dynamics of the process, yet our knowledge is incomplete. That is, we may not know the parameters of the mechanistic model and/or the mechanistic model is an incomplete representation of the process. An incomplete representation of dynamics could be due to many circumstances – for instance, due to approximations to the underlying partial differential equations, or in other instances due to missing mechanistic understanding. Next, we outline the notation to be used in this exposition.
@@ 2.1 Process and Data Notation
Assume we have observations of elements of the process corresponding to some or all of the times in $\mathcal{T}$ at $m_{t}$ spatial locations, which we denote by the $m_{t}$-vector $\bm{z}_{t}$. Some or all of these observation locations may coincide with the spatial locations in $\mathcal{S}$, but may not, as discussed below. Note, it may be the case that $m_{t}=0$ for some times (i.e., there are no observations for certain times).
@@ 2.2 Mechanistic Model Notation
Here we focus our attention on mechanistic models that take the form of a time-dependent partial differential equation (PDE), but the ideas presented here can be easily adapted to other forms of mechanistic models (e.g., ordinary differential equations, or agent-based models). Assume we have a PDE that we believe is mechanistically informative to our latent process of interest, $\bm{u}_{t}(\bm{s})$. We denote this as
$\frac{\partial\bm{v}_{t}(\bm{s})}{\partial{t}}=\mathcal{M}(\bm{v}_{t}(\bm{s}),\bm{\theta}_{m}),$ (1)
where $\bm{v}_{t}(\bm{s})$ corresponds to the $K$-dimensional mechanistic process defined at $t\in[0,T+\tau]$ and $\bm{s}\in\mathcal{D}$ and $\mathcal{M}$ is an operator controlled by parameters $\bm{\theta}_{m}$. For example, we may believe that our latent process shows evidence of advective and diffusive dynamics, both of which can be represented through appropriate terms in $\mathcal{M}$. Note, we are deliberately making a distinction between the true latent process of interest, $\bm{u}_{t}(\bm{s})$, and the deterministic mechanistic model representation, denoted by ${\bm{v}}_{t}(\bm{s})$ - this distinction will be made more clear when we discuss models in Section 3 and 4. We also assume that the PDE has associated initial and boundary conditions, respectively,
$\bm{v}_{0}(\bm{s})$ $=\bm{g}_{I}(\bm{s}),\;\;\bm{s}\in\mathcal{D},$
$\bm{B}(\bm{v}_{t}(\bm{s}))$ $=0,\;\;t\in[0,T+\tau],\bm{s}\in\mathcal{D}_{B},$
@@
Our goal is to perform inference on $\bm{u}_{t}(\bm{s})$ at the $\mathcal{S}$ spatial locations and at times $t\in\mathcal{T}$ while taking into account the information present in $\mathcal{M}$ and given the data $\bm{Z}\equiv\{\bm{z}_{1},\ldots,\bm{z}_{T}\}$. We can denote this as a posterior distribution (with some abuse of notation) as $\bm{u}_{t}:t\in\mathcal{T}|\bm{Z},\mathcal{M}]$, where we use the bracket notation ([][\;\;]) to represent a probability distribution, as is common in the Bayesian hierarchical modeling literature (e.g., gelfand1990sampling; cressie2011statistics); that is, $[a|b]$ corresponds to the distribution of $a$ given (conditional upon) $b$, $[b]$ corresponds to the marginal distribution of $b$, and $[a,b]$ corresponds to the joint distribution of $a$ and $b$. Typically, we are interested in learning the latent process at some initial time $\bm{u}_{0}$, interpolation at some time $t\in(0,T]$, $\bm{u}_{t}$, or forecasting at some time $T+\tau$, $\bm{u}_{T+\tau}$. This is generally termed the “forward problem” where $\mathcal{D},\mathcal{D}_{B},\bm{g}_{I}(\cdot),\bm{B}(\cdot)$, and $\bm{\theta}_{m}$ are known and one seeks to find $\bm{u}_{0}(\bm{s})$ and $\bm{u}_{t}]_{t\in(0,T+\tau]}$ given the dynamics suggested by $\mathcal{M}$. We may also be interested in learning parameters associated with the mechanistic PDE ($\bm{\theta}_{m}$) and various parameters associated with the underlying statistical or neural models used to represent the PDE. This is generally termed the “inverse problem,” where we may have partial observations of the process in the domain of interest. The primary goal of the methodology presented in this work is to infer the latent process $\{\bm{u}_{t}\}_{t\in[0,T+\tau]}$ on $\mathcal{D}$ as well as the $\bm{\theta}_{m}$ that satisfies $\mathcal{M}$ given observations $\bm{Z}$.
@@
$\bm{v}_{t+1}=\mathcal{M}(\bm{v}_{t};\bm{\theta}_{m}),$ (2)
but recognize that there are PDE solvers that provide $\bm{v}_{t}$ for any and all times in $[0,T]$ that are not necessarily sequential in time (e.g., implicit solvers).
@@ 3 Historical Perspective
3.1 Variational Data Assimilation
@@
It was known shortly after L.F. Richardson’s failed attempt at the first numerical weather forecast in the early 1920s (richardson1922weather) that the numerical solution of PDEs with initial conditions given by data would quickly become unstable due to lack of consistency with the governing equations. Thus, by the time that computers were actually capable of generating a realistic forecast in real time, it was known that initial condition fields of weather variables had to be smoothed in a way that respected dynamical constraints (charney1950numerical; phillips1956general). Two main approaches were developed in the 1950s and 1960s to do this, one deterministic based on the calculus of variations and one stochastic, based on what would come to be known as Gaussian processes (GPs). In this section, we briefly discuss these approaches as they provide the foundation for mechanistically-informed models for spatio-temporal processes in statistics.
@@
The initial idea in data assimilation, that one could minimize an objective function containing the sum of squared differences between the observations and the “true” initial field while simultaneously satisfying a constraint that incorporated the PDE dynamics, was presented in sasaki1958objective. Although he formulated this in continuous space, practical uses of the methodology considered discrete space (e.g., thompson1969reduction; sasaki1970some; stephens1970variational).
@@
Following the notation in Section 2.1, say we have $m_{t}$-dimensional data vectors, $\bm{z}_{t}$, and numerical solutions of the deterministic PDE state process at time $t$ denoted by $\bm{v}_{t}$ given by (2). Here, without loss of generality, we assume $K=1$ (a univariate process) for notational simplicity. Critically, for deterministic systems, if one knows the initial state $\bm{v}_{0}$ then (2) determines the sequence of future states, $\{{\bm{v}}_{1},{\bm{v}}_{2},\ldots,{\bm{v}}_{t},\ldots\}$. In Sasaki’s variational approach, one seeks an initial state ${\bm{v}}_{0}$ that combines the physics from the PDE and observations to obtain the initial condition. Importantly, in the context of weather forecasting, it is this estimated initial condition that is used to generate forecasts from the deterministic model, $\{{\bm{v}}_{t},t=1,\ldots\}$.
Sasaski’s approach estimates the initial state, $\bm{v}_{0}$, by minimizing an objective function, $J_{0}(\bm{v}_{0})$, subject to the constraint that the state evolution follows the numerical PDE solver exactly. That is, (e.g., see lewis2008sasaki) the optimization problem is to minimize
$J_{0}(\bm{v}_{0})=\frac{1}{2}\sum_{t=1}^{T}(\bm{z}_{t}-\bm{h}(\bm{v}_{t}))^{\prime}\bm{R}_{t}^{-1}(\bm{z}_{t}-\bm{h}(\bm{v}_{t})),$ (3)
@@
subject to the constraint of (2) – i.e., $\bm{v}_{t+1}=\mathcal{M}(\bm{v}_{t};\bm{\theta}_{m})$ for t=0,,T1t=0,...,T-1; this constraint allows each of $\bm{v}_{1},...,\bm{v}_{T}$ to be written as a function of $\bm{v}_{0}$ by iterating application of $\mathcal{M}(.;\bm{\theta}_{m})$. Additionally, $\bm{h}(\cdot)$ is a (known) mapping function that takes the $n$-dimensional deterministic model state vector $\bm{v}_{t}$ and maps to the $m_{t}$-dimensional observation vector at time $t$ (e.g., an interpolator if data are not coincident with model locations, or perhaps the identity if they are; this function can also take into account nonlinear transformations, see daley1993atmospheric). In addition, $\bm{R}_{t}$ is the time-varying $m_{t}\times m_{t}$ observation error covariance matrix (representing measurement error and other sources of observational error), also typically assumed known. lewis2008sasaki also describe how this constrained optimization problem can be solved via the method of Lagrange multipliers. We emphasize that this approach is considered a “strong” constraint because the state solution must follow the deterministic model exactly (sasaki1970some).